How long will it take a 50W heater to melt 100g of ice at 0 celsius degree
1
Expert's answer
2021-03-23T11:39:14-0400

"Q=Lm"


Q is the amount of heat needed to melt the ice 


L is the latent heat of melt for ice, that is


m  is the mass of ice that is to melt


"m=100g=0.1kg"


"Q=3.34.10^{5}\times0.1= 33.4kJ"


We know that "Power=\frac {E}{P}"


Therefore"t=\frac{E}{p}= \frac{33.4kJ}{50}"


"t=668 seconds"